OpenStudy (anonymous):

x^2y^2=cos(x^2+y^2) solve for y-prime :)

jimthompson5910 (jim_thompson5910):

x^2y^2 = cos(x^2+y^2) d/dx(x^2y^2) = d/dx(cos(x^2+y^2)) 2xy^2 + 2x^2y*y' = -sin(x^2+y^2)*(2x+2y*y') 2xy^2 + 2x^2y*y' = -2x*sin(x^2+y^2)-2y*y'*sin(x^2+y^2) 2x^2y*y'+2y*y'*sin(x^2+y^2) = -2x*sin(x^2+y^2)-2xy^2 y'(2x^2y+2y*sin(x^2+y^2)) = -2x*sin(x^2+y^2)-2xy^2 y' = (-2x*sin(x^2+y^2)-2xy^2)/(2x^2y+2y*sin(x^2+y^2))

OpenStudy (anonymous):

@jim_thompson5910 what happened here:-2x*sin(x^2+y^2)-2y*y'*sin(x^2+y^2) the 3rd line thats where i got stocked

jimthompson5910 (jim_thompson5910):

On the left side, i use the product rule. On the right side, I used the chain rule.

jimthompson5910 (jim_thompson5910):

oh and wherever you're differentiating y wrt x, then you're using the chain rule as well

OpenStudy (anonymous):

oh ok! thanks alot :)

jimthompson5910 (jim_thompson5910):

you're welcome